# Robustness Optimization for Compact Free-electron Laser Driven by Laser Wakefield Accelerators

### v4

@@ Abstract
Figure 5: Radiation energy as a function of beam pointing jitter. Mean values (black squares) and standard deviation (red error bars) derived from 20 independent simulations (red dots) with random seeds.
@@
In conclusion, we present a systematic design and optimization of compact LWFA-driven FELs with enhanced robustness and reliability. Through comprehensive start-to-end simulations, the feasibility of the proposed scheme has been demonstrated, achieving pulse energies exceeding 1 μJ under realistic experimental conditions. These simulations account for inherent parameter fluctuations, including laser energy variation, focal position displacement induced by wavefront distortion, and shock front position instability in the plasma. The optimized configuration exhibits remarkable tolerance to these critical parameter fluctuations while maintaining optimal phase space matching between the LWFA electron bunches and the undulator radiation field. These results establish crucial technical foundations for developing table-top FELs with unprecedented compactness, particularly highlighting robustness of the system against parameter jitter characteristic of current experimental configurations of LWFAs. To the best of our knowledge, the proposed scheme represents the first demonstration of μJ-level FEL radiation that is compatible with the stability issues of LWFA under actual experimental conditions.
@@ Acknowledgements.
This work was supported by National Natural Science Foundation of China (Grant Nos. 12388102, 12225411, 12474349, 12105353, 12174410 and 12434012), Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0890201), CAS Project for Young Scientists in Basic Research (Grant No. YSBR060), Program of Shanghai Academic Research Leader (Grant No. 22XD1424200), the State Key Laboratory Program of the Chinese Ministry of Science and Technology and CAS Youth Innovation Promotion Association (2022242).
Appendix A Appendix A: Measurement of shock front
The shock front is essential for the injection process in LWFA and thus significantly impacts the beam quality. As previously discussed, the position jitter with an RMS value of 4.9 μm was used in our simulations, which was experimentally measured. To capture the shock front properties, a probe laser passed through the plasma region and imaged using a 4f imaging system before entering the interferometer. The resulting interference patterns were captured using a charged-coupled device (CCD), with the shock front region depicted in Fig. 6(a). The corresponding phase distribution and the first-order derivative of the fringe phase are illustrated in Figs. 6(b) and  6(c), respectively. To characterize the position jitter of the shock front, we performed continuous acquisition of 100 experimental shots, as shown in Fig. 6(d). An RMS value of $\sigma_{zshock}=4.9$ μm was obtained. The statistical parameter is represented in the red shaded region in Fig. 6(d), providing quantitative analysis of the shock wave stability.
Figure 6: Properties of the measured shock front. (a) Measured interferometric fringe patterns and (b)-(c) the corresponding distributions of phase (b) and its first-order derivative (c). (d) Statics on the shock front jitter over consecutive 100 shots based on the first derivative of phase. The red and gray shaded region denotes the RMS and ±$\pm 2$ times RMS range, respectively.
Appendix B Appendix B: Gain length optimization
@@
$\eta_{\epsilon}$ $=$ $\frac{L_{G0}}{2k_{r}\sqrt{\beta_{x}\beta_{y}}\sqrt{\epsilon_{x}\epsilon_{y}}}\;,$ (4)
$\eta_{\gamma}$ $=$ $\frac{L_{G0}}{4\pi\lambda_{u}\sigma_{\eta}}\;.$ (5)
@@
where $\gamma_{0}$ is the electron energy in units of the rest energy, $I_{p}$ is the peak current, $I_{A}=17$ kA is the Alfvén current, $K_{0}$ is the undulator parameter, $[JJ]=[J_{0}(\xi)-J_{1}(\xi)]$ with $\xi=K_{0}^{2}/(4+2K_{0}^{2})$ is the Bessel function factor for a planar undulator, $K_{u}=2\pi/\lambda_{u}$, $k_{r}=2\pi/\lambda_{r}$, $\lambda_{r}=\lambda_{u}(1+K_{0}^{2}/2)/(2\gamma_{0}^{2})$ is radiation wavelength, $\sigma_{x}$ and $\sigma_{y}$ is the average beam size in the undulator, $\beta_{x}$ and $\beta_{y}$ is the average Twiss parameter in the undulator, $\epsilon_{x}$ and $\epsilon_{y}$ is the average emittance in the undulator, and $\sigma_{\eta}$ is the relative RMS slice energy spread. The parameters $\gamma_{0}$, $I_{p}$, and $\sigma_{\eta}$ were obtained from FBPIC numerical simulations, whereas the transverse beam characteristics $\sigma_{x}$, $\sigma_{y}$, $\beta_{x}$, $\beta_{y}$, $\epsilon_{x}$, and $\epsilon_{y}$ were derived from particle tracking simulations using the ELEGANT code [37].
@@
In the optimization framework targeting $<L_{G}^{-1}>$, we conducted parallel particle tracking simulations utilizing the 13-electron-beam ensemble illustrated in Fig. 1. The resulting transverse beam characteristics were subsequently employed to calculate $<L_{G}^{-1}>$, which serves as the primary quantitative metric for objective function evaluation. The parameter ranges for the optimization are presented in Table 2.
Table 2: Parameter space used in $<L_{G}^{-1}>$
@@
optimization. Parameter Min. value Max. value Best value [^2] Note $l_{0}-l_{6}$ in units of meters (m), $q_{1}-q_{6}$ in units of per square meter (m-2). $l_{0}$ 0.03 1.0 0.08 $l_{1}$ 0.03 1.0 0.11 $l_{2}$ 0.03 1.0 0.16 $l_{3}$ 0.3 3.0 1.17 $l_{4}$ 0.3 3.0 2.98 $l_{5}$ 0.3 3.0 0.39 $l_{6}$ 0.3 3.0 0.33 $q_{1}$143 143 -139 $q_{2}$143 143 141 $q_{3}$143 143 -69 $q_{4}$45 45 -0.34 $q_{5}$45 45 10.51 $q_{6}$45 45 -10.82
@@
Figure 7: The optimization results of the $<L_{G}^{-1}>$. (a) The simulated objective function $<L_{G}^{-1}>$ (dots) with the cumulative best results (black dashed line). (b)-(c) The evolution of the beam size (b) and the normalized projected emittance (c) along the beamline in horizontal (blue) and vertical (orange) directions. The tracking beamline was configured using the best parameters obtained from the optimization.
@@
Figure 7(a) illustrates the evolution of the objective function $<L_{G}^{-1}>$ throughout the BO process, with the cumulative optimal results represented by the black dashed line. The optimization commenced with random sampling within the predefined parameter space detailed in Table 2. A progressive convergence of the objective function emerges after approximately 1200 iterations, with the optimal solution achieved at iteration 10,071, corresponding to an average 3D power gain length of approximately 0.24 m. The resulting beam sizes and transverse emittance for the 13-electron-beam ensemble under the optimized conditions are presented in Figs. 7(b) and  7(c), respectively. It should be noted that slice-averaged energy spread of the e beam is employed when calculating the power gain length. The radiation gain predominantly originates from specific slices with advanced beam qualities, while other slices contribute minimally to the overall gain. The optimization based on slice-averaged values may cause potential deviations, considering the different properties over slices of the e beam from LWFAs. To address this limitation, the optimization using the radiation energy as an alternative objective function is implemented.

### v5

@@ III BEAMLINE OPTIMIZATION
Figure 4: Radiation energy as a function of (a) normalized vector potential $\Delta a_{0}$, (b) laser focal position $\Delta{z_{foc}}$ and (c) the shock-front position $\Delta{z_{shock}}$. Mean values (black squares) and standard deviation (error bars) calculated from 20 independent simulations with varying random seeds, where the blue curve represents optimization using $L_{G}^{-1}>$ as objective function and the red curve represents optimization using $E_{min}$ as the objective function. The y-axis is plotted on a logarithmic scale.
Figure 3(a) illustrates the evolution of the objective function $E_{min}$ as a function of iterations during the optimization. It is noted that the radiation energy for each evaluation was derived in a single simulation run using a fixed random seed for speed up. The optimal solution achieves at iteration 11,020, with the corresponding optimal parameters detailed in Table 1. Under these optimized conditions, the radiation energy consistently exceeds 1 µJ for all electron beams in the ensemble, indicating robust operation of the LWFA-driven FELs in the saturated or high-gain regime(as shown in Fig. 5(a)) and effectively accommodating shot-to-shot fluctuations. The corresponding evolution of the beam sizes and transverse emittance along the beamline is presented in Figs. 3(b) and  3(c), respectively. Figures 4(a)- 4(c) provide the comparison of the two optimization frameworks targeting $L_{G}^{-1}>$ and $E_{min}$ under their respective optimized conditions. It is noted that the FEL energy was simulated with distinct random seeds for 20 runs in each situation in Fig. 4. While both frameworks exhibit similar trends in radiation energy, the $E_{min}$ optimized framework demonstrates superior performance with higher radiation energy. This enhancement can be attributed to two aspects, the fitting distortion of Eq. (B) for the short-pulse situation induced by the slippage effects [35], and the potential deviations introduced by using slice-averaged parameters for power gain length calculations. Through numerical simulations, the slice beam properties can be optimized to simultaneously achieve localized small averaged beam size, low energy spread, and high beam current, where the lasing occurs. Such optimization ensures more efficient electron-radiation interaction and facilitates energy extraction efficiency from the e beam.
Figure 5: Simulation results for FELs with the 13-electron-beam ensemble. (a) Radiation energy along the undulator with the y-axis on a logarithmic scale. (b) Pulse duration and (c) spectrum of the generated radiation pulses at the exit of the undulator. Each line represents the average value of 20 distinct simulations with random seeds.
Figure 5 presents the properties of the FEL radiation under the optimal operation targeting $E_{min}$. The evolution of the radiation along the undulator for the 13-electron-beam ensemble is shown in Fig. 5(a). A small subset of the beams enters the exponential-gain regime at about 1 m indide the undulator, and the majority begin exponential growth beyond 1.5 m and approach saturation or near-saturation by the undulator exit. Figure 5(b) illustrates the radiation power, highlighting localized lasing enabled by the carefully designed beamline. Through energy spread minimization and emittance matching, the beam parameters locally improve the phase-space characteristics, thereby enhancing the gain dynamics. Notably, even the minimum radiation power remains at the sub-gigawatt level, demonstrating strong tolerance to system fluctuations. The corresponding radiation wavelength fluctuates with an RMS value of 0.8 nm, as shown in Fig. 5(c) and Table 2. This variation originates from energy fluctuations of the accelerated e-beam, and thus the resonant wavelength. Detailed FEL simulation results for all 13 beam ensembles are summarized in Table 2. It should be noted that the results presented in Fig. 5 and Table 2 correspond to average values obtained from 20 independent simulations, each initialized with a random seed.
Table 2: Simulation results for FELs with the 13-electron-beam ensemble.
Parameter jitter Energy (µJ) Peak power (GW) Pulse duration (fs) [^2] Note that the pulse duration is defined as the full width at half maximum (FWHM). Wavelength (nm)
$2\sigma_{a_{0}}$ 2.09 1.26 1.40 24.35
$\sigma_{a_{0}}$ 6.06 3.65 1.45 24.58
$\sigma_{a_{0}}$ 27.51 13.27 1.75 25.04
$2\sigma_{a_{0}}$ 4.91 1.54 2.45 25.77
$2\sigma_{zfoc}$ 1.58 0.31 5.16 23.91
$\sigma_{zfoc}$ 8.10 2.79 2.85 24.73
$\sigma_{zfoc}$ 35.84 17.18 1.78 24.96
$2\sigma_{zfoc}$ 11.35 5.32 1.90 25.12
$2\sigma_{zshock}$ 4.28 1.05 3.87 23.22
$\sigma_{zshock}$ 6.44 3.17 1.48 24.28
$\sigma_{zshock}$ 20.10 8.63 2.03 25.36
$2\sigma_{zshock}$ 2.93 0.77 3.22 26.54
jitter-free 33.61 11.15 3.19 25.52
IV Discussion and Conclusions
Figure 6: Parameters of the optimized electron beam: (a) power gain length, (b) Pierce parameter, and (c) energy extraction efficiency. The blue, red and orange curves represent the laser energy variations, laser focal position displacement and shock front position instability, respectively. Horizontal axes denotes $\pm 2$ folds RMS range of these three parameter jitters.
Figures. 6(a) and  6(b) show the 3D gain length and the Pierce parameter under the optimized condition, and the corresponding calculation is detailed in Appendix B. Figure 6(c) presents the FEL energy-extraction efficiency, defined as the ratio of the peak radiation power $P_{r}$ to the peak electron-beam power $P_{e}=I\cdot E/e$, where $I$ is the peak current, $E$ is the beam energy, and $e$ is the elementary charge. The saturation length in a SASE FEL, defined as the undulator length at which the radiation power saturates, scales with the power gain length $L_{G}$. The scaling factor exhibits a logarithmic dependence on the number of electrons within one coherence time and typically lies within a narrow range of 18 to 20 [21]. When accounting for specific parameter jitters in Fig. 6(a), the predicted saturation length exceeds 4.5 m. This theoretical prediction aligns with the results presented in Fig. 5(a), where the saturation is not reached within the considered undulator length for some cases. Extending the undulator would allow the radiation to progress from exponential gain regime to full saturation, thereby fulfilling its potential for higher energy output.
Beam pointing stability is another critical issue that must be addressed to ensure the robust operation of LWFA-driven FELs. To quantitatively evaluate the pointing instability, artificial angular kicks were introduced at the beamline entrance using e beams simulated under jitter-free conditions ($\Delta_{RMS}=0$). Angular deviations in both transverse planes ($x^{\prime}:0.1-1.0$ mrad, $y^{\prime}:0.1-1.0$ mrad) were systematically introduced. Figure 7 quantifies the resulting radiation energy degradation caused by this angular perturbation, revealing a monotonic decline in radiation energy from 34 µJ to 1 µJ as the pointing deviation increases to 1 mrad. Notably, the average radiation energy over 20 random seeds remains above 1 µJ even with 1 mrad deviation, indicating the proposed scheme’s remarkable tolerance to angular deviations.
Figure 7: Radiation energy as a function of beam pointing jitter. Mean values (black squares) and standard deviation (red error bars) derived from 20 independent simulations (red dots) with random seeds. The y-axis is plotted on a logarithmic scale.
@@ IV Discussion and Conclusions
In conclusion, we present a systematic design and optimization of compact LWFA-driven FELs with enhanced robustness and reliability. The feasibility of the proposed scheme is demonstrated through comprehensive start-to-end simulations, which consistently yield FEL energies exceeding 1 µJ, even when accounting for laser and plasma instabilities. These simulations explicitly incorporate inherent parameter fluctuations, including laser energy jitter, focal position displacement induced by wavefront distortion, and shock front position instability in the plasma. The optimized configuration exhibits remarkable tolerance to these critical parameter fluctuations while maintaining optimal matching between the LWFA e beams and the radiation field. These results establish crucial technical foundations for developing table-top FELs with unprecedented compactness, particularly highlighting robustness of the system against parameter jitter characteristic of current experimental configurations of LWFAs.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grant Nos. 12388102), the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0890200), the National Natural Science Foundation of China (12225411, 12474349, 12174410 and 12434012), CAS Project for Young Scientists in Basic Research (Grant No. YSBR060), CAS Youth Innovation Promotion Association (No. 2022242), and the New Cornerstone Science Foundation through the XPLORER PRIZE.
Data availability
The data that support the findings of this article are openly available [26].
Appendix A Measurement of shock front
The shock front is essential for the injection process in LWFA and thus significantly impacts the beam quality. As previously discussed, the position jitter with an RMS value of 4.9 µm was used in our simulations, which was experimentally measured. Such a density profile can be achieved via the interaction of supersonically expanding helium gas with baffle. To capture the shock front properties, a probe laser passed through the plasma region and imaged using a 4f imaging system before entering the interferometer. The resulting interference patterns were captured using a charged-coupled device (CCD), with the shock front region depicted in Fig. 8(a). The corresponding phase distribution and the first-order derivative of the fringe phase are illustrated in Figs. 8(b) and  8(c), respectively. To characterize the position jitter of the shock front, we performed continuous acquisition of 100 experimental shots, as shown in Fig. 8(d). An RMS value of $\sigma_{zshock}=4.9$ µm was obtained. The statistical parameter is represented in the red shaded region in Fig. 8(d), providing quantitative analysis of the shock wave stability.
Figure 8: Properties of the measured shock front. (a) Measured interferometric fringe patterns and (b)-(c) the corresponding distributions of phase (b) and its first-order derivative (c). (d) Statics on the shock front jitter over consecutive 100 shots based on the first derivative of phase. The red and gray shaded region denotes the RMS and $\pm 2$ times RMS range, respectively.
Appendix B Gain length optimization
@@
$\eta_{\epsilon}$ $=$ $\frac{2L_{G0}k_{r}}{\sqrt{\beta_{x}\beta_{y}}}\sqrt{\epsilon_{x}\epsilon_{y}}\;,$ (4)
$\eta_{\gamma}$ $=$ $\frac{4\pi L_{G0}}{\lambda_{u}}\sigma_{\eta}\;.$ (5)
@@
where $\gamma_{0}$ is the electron energy in units of the rest energy, $I_{p}$ is the peak current, $I_{A}=17$ kA is the Alfvén current, $K_{0}$ is the undulator parameter, $JJ]=[J_{0}(\xi)-J_{1}(\xi)]$ with $\xi=K_{0}^{2}/(4+2K_{0}^{2})$ is the Bessel function factor for a planar undulator, $K_{u}=2\pi/\lambda_{u}$, $k_{r}=2\pi/\lambda_{r}$, $\lambda_{r}=\lambda_{u}(1+K_{0}^{2}/2)/(2\gamma_{0}^{2})$ is radiation wavelength, $\sigma_{x}$ and $\sigma_{y}$ is the average beam size in the undulator, $\beta_{x}$ and $\beta_{y}$ is the average Twiss parameter in the undulator, $\epsilon_{x}$ and $\epsilon_{y}$ is the average emittance in the undulator, and $\sigma_{\eta}$ is the relative RMS slice energy spread. The parameters $\gamma_{0}$, $I_{p}$, and $\sigma_{\eta}$ were obtained from FBPIC numerical simulations, whereas the transverse beam characteristics $\sigma_{x}$, $\sigma_{y}$, $\beta_{x}$, $\beta_{y}$, $\epsilon_{x}$, and $\epsilon_{y}$ were derived from particle tracking simulations using the ELEGANT code [3].
@@
In the optimization framework targeting $L_{G}^{-1}>$, we conducted parallel particle tracking simulations utilizing the 13-electron-beam ensemble illustrated in Fig. 1. The resulting transverse beam characteristics were subsequently employed to calculate $L_{G}^{-1}>$, which serves as the primary quantitative metric for objective function evaluation. The parameter ranges for the optimization are presented in Table 3.
Table 3: Parameter space used in $L_{G}^{-1}>$
@@
optimization. Parameter Min. value Max. value Best value [^3] Note $l_{0}-l_{6}$ in units of meters (m), $q_{1}-q_{6}$ in units of per square meter (m-2). $l_{0}$ 0.03 1.0 0.08 $l_{1}$ 0.03 1.0 0.11 $l_{2}$ 0.03 1.0 0.16 $l_{3}$ 0.3 3.0 1.17 $l_{4}$ 0.3 3.0 2.98 $l_{5}$ 0.3 3.0 0.39 $l_{6}$ 0.3 3.0 0.33 $q_{1}$ -143 143 -139 $q_{2}$ -143 143 141 $q_{3}$ -143 143 -69 $q_{4}$ -45 45 -0.34 $q_{5}$ -45 45 10.51 $q_{6}$ -45 45 -10.82
@@
Figure 9: The optimization results of the $L_{G}^{-1}>$. (a) The simulated objective function $L_{G}^{-1}>$ (dots) with the cumulative best results (black dashed line). The y-axis is plotted on a logarithmic scale. (b)-(c) The evolution of the beam size (b) and the normalized projected emittance (c) along the beamline in horizontal (blue) and vertical (orange) directions. The tracking beamline was configured using the best parameters obtained from the optimization.
@@
Figure 9(a) illustrates the evolution of the objective function $L_{G}^{-1}>$ throughout the BO process, with the cumulative optimal results represented by the black dashed line. The optimization commenced with random sampling within the predefined parameter space detailed in Table 3. A progressive convergence of the objective function emerges after approximately 1200 iterations, with the optimal solution achieved at iteration 10,071, corresponding to an average 3D power gain length of approximately 0.24 m. The resulting beam sizes and transverse emittance for the 13-electron-beam ensemble under the optimized conditions are presented in Figs. 9(b) and  9(c), respectively. It should be noted that slice-averaged energy spread of the e beam is employed when calculating the power gain length. The radiation gain predominantly originates from specific slices with advanced beam qualities, while other slices contribute minimally to the overall gain. The optimization based on slice-averaged values may cause potential deviations, considering the different properties over slices of the e beam from LWFAs. To address this limitation, the optimization using the radiation energy as an alternative objective function is implemented.